OpenStudy (anonymous):

Darla is building a new desk. To make sure she had made a square corner, she measures 9ft from the corner along one edge and 7ft from the corner along the other edge. How long should the diagonal be between those two points if the corner is a right angle? Diagonal ? ft

OpenStudy (anonymous):

\[\sqrt{7^2+9^2}=\sqrt{49+81}=\sqrt{130}\approx11.40\]

OpenStudy (aravindg):

11.40